Slope of 2/9 passes through (-3,2) in slope intercept form and standard
Answer by mananth(16946)

Slope of 2/9 passes through (-3,2) in slope intercept form and standard
The equation in point slope form = (y-y1)=m(x-x1)
y-2=2/9(x+3)
9(y-2) =2x+6
9y-18 =2x+6
9y= 2x+24
y = 2/9 x +24/9
y=(2/9)x+8/3